evaluate
square root of
4 over 25

Answer by checkley75(3666)

WHAT YOU ARE NOT DOING IS TAKING THE SQUARE ROOT OF BOTH NUMERATOR & DENOMINATOR. YOU MUST TAKE THE SQRT OF ALL VALUES UNDER THE RADICAL.
EXAMPLE: SQRT(16/64)=SQRT(4*4/8*8)=4/8=1/2 OR SQRT(16/64)=SQRT(1/4)=1/2
--------------------------------------------------------------------------------
SQRT(4/25)=SQRT(2*2/5*5)=2/5